# Explaining mountain pine beetle dynamics: From life history traits to large, episodic outbreaks

### v1

@@ 2.1 Overview
To understand the effects of MPB life-history traits on MPB population dynamics, we will build a mechanistic model of MPB dynamics (Fig. 6), then simulate the model under counterfactual scenarios where various life-history traits have been turned on or off. If a phenomenon (e.g., episodic outbreaks) disappears when a certain life-history trait is turned off, we can say that the life-history trait is a necessary condition for the phenomenon. If two traits being turned off makes the difference, then the two traits are jointly necessary. Of course, our inferences are conditional on nature following the dynamics described by the model.
Figure 6: Graphical representation of key model processes.
Realism is the key feature of our model — we intend to emulate the actual ecological system so that we can conduct virtual experiments. Many models of MPB have been published, but many are too simple for our purposes, having favored mathematical tractability over realism. Other models are too complex, containing functional forms that cannot be reasonably parameterized given available data. For example, while we understand that beetles are attracted to a suite of pheromones, explicitly modelling the production and diffusion of these pheromones — which depend on the stand structure, transient wind conditions, and the natal hosts of attacking female beetles — is a herculean task. Fitting such a model requires that we fix a large proportion of parameters (e.g., Biesinger et al.,, 2000, Table 2), yet we lack the appropriate kind of data to select these parameter values with confidence. More generally, the pursuit of realism in our model is tempered by the inherent complexity of the MPB system and data limitations.
In this section, our overarching principle is that all model structures be thoroughly justified via clear patterns in the data and/or a preponderance of evidence from previous research. Beyond the justifications outlined in the model description (below), additional support is provided in Appendix B. Our model, which examines the interactions between MPB and lodgepole pine over a one square mile area, is validated and parameterized using data from this spatial scale. Given that beetles primarily disperse short distances (refer to Section 1.1), inter-patch dispersal has minimal impact on local populations and is not explicitly modeled. Additionally, some processes related to within-stand dispersal are implicitly modelled. For example, the aforementioned diffusion of pheromones is subsumed under a linear relationship between beetle preference and tree diameter, since larger trees produce larger plumes of attractive terpenoids.
@@ 2.2 Data
The model was predominantly validated and parameterized using data from Boone et al., (2011) and Klein et al., (1978). All data was extracted from figures and tables using the GraphGrabber software and Microsoft Excel’s digitizing features. Specific variables are describe in Table A.1. The Boone et al., dataset contains the number of mass attacked trees and the number of successfully mass attacked trees, across a range of attack years, stands, and attack densities. This data justifies a functional form for the Allee effect. The second dataset, courtesy of Klein et al., (Fig. 7), is one of the only datasets that contain beetle attacks, beetle emergence, and tree deaths, all stratified by year and tree diameter class. Our statistical model of MPB outbreak dynamics is primarily calibrated with the Klein et al., dataset, with the notable exception being that a Bayesian prior for the Allee effect parameter is derived from the Boone et al., data (more details in Section 2.3).
Klein et al.,’s study was conducted from 1965–1972 in the southwest corner of Yellowstone National Park in Idaho. Data was collected and aggregated across 85 permanent plots that spanned a square mile. Trees were grouped into six 2-inch diameter classes, ranging from 8 to $\geq$ 16 inches. For each of the 106 sampled trees, paired 6x6-inch sample squares were marked on opposite sides of the bole at regular height intervals. Within these squares, emergence holes and successful attacks (operationalized as egg-gallery starts) were measured. Some of this data was originally presented by Parker, (1973), but a more thorough reporting is given by Klein et al., (1978).
Figure 7: Time series of emergence holes and gallery starts in the Klein et al., dataset. Emergence and subsequent attacks are assumed to take place in July and August respectively. There are always far fewer gallery starts than emergence holes, suggesting high levels of dispersal mortality.
2.3 Simulation model description
2.3.1 Outbreak dynamics
Throughout the model description, we will refer to the surface area over which beetles may attack as the susceptible surface area; note that beetles do not attack the upper part of the bole due to thin phloem. To convert between tree diameter and susceptible surface area, we use a simple linear model,
$SA=-3.792+0.821\times DBH,$ (1)
We start with beetle density in year $y$, denoted $B_{y}$ (units: females per acre). We do not track males, since previous research has shown that nearly all females are mated (Reid,, 1958; Amman,, 1980). Upon emergence, beetles disperse in search of a new host. The probability of survival during this dispersal phase is
$s_{y}=s_{0}\lambda_{y},$ (2)
where $s_{0}$ is the maximum survival probability, and $\lambda_{y}$ is a factor (between 0 and 1) that reduces the maximum survival probability when many trees have been killed, putatively because beetles are dispersing farther and thus suffering from higher dispersal mortality.
The effect of stand-dependent dispersal on the dispersal-phase survival is
$\lambda_{y}=\left(1+\exp\left(-\lambda_{1}\left((\text{relative surface area})_{y}-\lambda_{0}\right)\right)\right)^{-1}.$ (3)
The right-hand side is simply a logistically-increasing function of the independent variable, $(\text{relative surface area})_{y}$, which is defined as the susceptible surface area of all trees with $DBH\geq 8\,\text{inches}$ in year $y$, divided by the total susceptible surface area when MPB are absent and the forest is at equilibrium. We use the relative surface area mostly for interpretability but also for computational reasons (the Bayesian model-fitting program Stan prefers all parameters to be on the unit-scale). The parameter $\lambda_{0}$ is the threshold for dispersal (i.e., $\lambda_{y}<0.5$ when $(\text{relative surface area})_{y}<\lambda_{0}$), and $\lambda_{1}$ controls the slope of the logistic function.
We cannot directly justify the functional form of Eq. 3 — say, by plotting total emergence against subsequent total attacks — because Klein et al., only reported attacks on successfully killed trees. However, we note that the logistic function is a standard way to model probabilities, and that the extreme limit of the function (i.e., dispersal mortality is near 100% when there are few trees) is supported by both the Klein et al., dataset (Fig. D.1) and the literature (see Section 1.2.2). In particular, Powell and Bentz, (2014) estimated that mean dispersal distance increased from several meters to several kilometers as tree density decreased.
@@ 2.3.1 Outbreak dynamics
To obtain the density of attacking beetles, we multiply the emerging females’ density and the dispersal-phase survival probability. Potentially, with probability $P_{I}$, we also add $I$ beetles per acre:
@@
The Bernoulli random variable simulates the stochastic endemic $\rightarrow$ incipient-epidemic transition. Because most MPB outbreaks occur during continent-scale epidemics, dispersal is the proximal catalyst for the endemic $\rightarrow$ incipient-epidemic transition on the local scale. But because drought is responsible for region-wide outbreaks, drought is the ultimate catalyst for most local-scale outbreaks. Drought conditions are somewhat unpredictable, but the best predictor of drought in western North America is the Pacific Decadal Oscillation (McCabe et al.,, 2004; Hidalgo,, 2004), which has a quasi-period of 15-30 years (Mantua and Hare,, 2002). The last two drought periods preceded the last two continent-scale outbreaks and were 20 years apart, (Creeden et al.,, 2014, Fig. 5d). Therefore, we set $P_{I}=1/20=0.05$.
To simulate the population size corresponding to the incipient-epidemic phase, we set $I=300$. This number was obtained with a back-of-the-envelope calculation: the median emergence density of females across several studies and Forest Service reports (Other Studies in Fig. 16), multiplied by the susceptible surface area of typical killed tree (10 inches DBH as an input to Eq. 1), multiplied by the median estimate of the dispersal-phase survival probability (using gallery-starts/emergence-holes quotient from the Klein et al., dataset, Fig. D.1), multiplied by 0.81, which is the number of successfully attacked trees per acre that corresponds to an incipient-epidemic state (according to Carroll et al., 2006c, ). The chosen values of $P_{I}$ and $I$ are educated guesses, but we re-ran our analysis with several different values and obtained qualitatively identical results.
After the dispersal phase, the remaining beetles attempt to mass attack pine trees, with the primary determinant of success being the number of nearby beetles. If beetles are unsuccessful, the tree survives and the beetles are assumed to die. This is a simplifying assumption because unsuccessful beetles could attack other trees; however, it is a decent assumption because some beetles will be killed by the first tree, the remainder face increased mortality risk during secondary dispersal, and late-arriving beetles have low fitness (Raffa,, 2001).
@@
where $a$ is the Allee threshold parameter. More precisely, $a$ is the beetle attacks per acre where the probability of success equals $0.9$. Thus, the number 2.3 in the above equation is obtained by solving $0.9=1-\exp(-x)$ for $x$.
The exponential saturating function in Eq. 5 is justified by the Boone et al., dataset (Fig. 8). A notable discrepancy is that the maximum probability of success reaches (close to) 100% in the model, but not in reality. However, an additional parameter for the maximum probability of successful attacks would be non-separable with the parameter $s_{0}$ (in Eq. 2). Thus, our estimate of $s_{0}$ not only represents maximum survival during the dispersal phase, but also the fact that some mass attacks are unsuccessful despite high beetle density. Additionally, $s_{0}$ captures low-vigor females that successfully attack trees but do not oviposit (5–30% of females could fall into this category; Amman,, 1975, 1980).
Figure 8: The probability of an attacked tree being killed is an increasing function of beetle density. The data is effectively modelled by an exponential saturating function (see Eq. 5, “Probability of success” term). The prediction was generated using posterior means of model parameters. Data from the Boone et al.,.
Next, the successfully attacking beetles distribute themselves between different tree diameter classes. To visualize the extent of MPB’s attack preference, we can plot the total beetle attacks across all trees within a diameter class, as a function of the proportion of tree surface area belonging to that diameter class (Fig. 9). If total beetle attacks are proportional to total tree surface area, then observations should cluster around the one-to-one line. Instead, we see that large trees are favored and small trees are disfavored.
Figure 9: Beetles disproportionately kill large trees. The x-axis is the proportion of susceptible surface area belonging to live trees of a particular diameter class, within a year. The y-axis is the proportion of total attacks received by the same diameter class, within a year. The black line is a LOESS model prediction. The grey dashed line is the one-to-one line, representing the no-mass-attack-preference hypothesis. Data from Klein et al., (1978).
Figure 10: Beetle preference is an increasing function of tree diameter. The data (from Klein et al.,, 1978) is well-modelled by a piecewise linear function (see Eq. 6). There is clear heteroskedasticity in the data, which is likely due to smaller sample sizes within classes of large-diameter trees. The exact number of samples is not available, so we use the within-DBH-class surface area as a proxy for sample size, visualized as scatter-plot point size. The prediction was generated using posterior means of model parameters.
We can operationalize mass attack preference by dividing the proportion of the beetle attacks by the proportion of tree surface area; that is, by dividing the y-coordinate by the x-coordinate in Figure 9. Note that this quantity is equivalent to the relative attack density: the attack density within a particular diameter class, divided by the attack density across all trees (within a year). Mass attack preference has a clear linear relationship with tree diameter (Fig. 10), which justifies the piecewise linear function,
$(\text{preference})_{j}=\begin{cases}\beta_{0}+\beta_{1}\cdot\text{{DBH}}_{j}&\text{if }\text{{DBH}}_{j}>-\beta_{0}/\beta_{1}\\ 0&\text{otherwise},\end{cases}$ (6)
where $\beta_{0}$ and $\beta_{1}$ are the intercept and slope parameters respectively. The subscripts $j$ and $y$ denote the DBH class and the year index respectively. The quantity $-\beta_{0}/\beta_{1}$ is the susceptible diameter threshold, i.e., trees with $\text{DBH}<-\beta_{0}/\beta_{1}$ are not attacked.
[…]
The equation for successful attacks, Eq. 7, shows how beetles are distributed among tree diameter classes, but not how beetles are distributed among individual trees. As it turns out, killed trees receive an approximately invariant number of attacks per unit surface area. That is, attack density does not depend on tree diameter (Fig. 11, Reid,, 1963, Table 1, Peterman,, 1974, Figs. 4–7). There is less literature on the relationship between attack density and beetle population size, but any such relationship is likely weak, given that 1) Negrón, (2018) found no relationship; 2) there is a clear mechanism for limiting attack density (see the following paragraph), which explains 3) that attack density does not depend on tree diameter, and 4) that attack density varies in the range of 50-100 females per square meter, despite beetle population sizes varying by several orders of magnitude. In the Klein et al., dataset, the link between attack density and beetle population size is inconclusive, but this has a negligible effect in any case (Appendix B.2).
Figure 11: Successful attack density (gallery starts per square meter) increases slightly over time, but does not depend on tree diameter at breast height (DBH). Data from Klein et al., (1978).
The mean attack density of 71 females per square meter in the Klein et al., dataset is close to the optimal value of 60 females per square meter, proposed by Raffa and Berryman, (1983). The mechanism for attaining near-optimal attack density is a suite of aggregation and anti-aggregation pheromones, and the relationship between such pheromones and host tree defenses. Beetles are attracted to the combination of terpenoid aggregation pheromones and volatilized terpenes. Because both are derived from tree resin, trees become less attractive as their resin defenses are depleted. Anti-aggregation pheromones, which are only detected by beetles at short distances (around a meter; Bentz,, 1996), redirect beetles to nearby trees, while rivalry stridulations (noises produced by males) help ensure that galleries are evenly spaced. Raffa, (2001) puts forward several hypotheses for the apparent lack of cheaters — beetles who ignore the anti-aggregation pheromones and enter a “full” tree.
Because of the invariant attack density, the number of killed trees per acre in the $j^{\text{th}}$ diameter class, denoted by $(\text{KTPA})_{j,y}$, is given by a linear function
@@
$(\text{KTPA})_{j,y}=\text{min}\left(\gamma\frac{(\text{succ. attacks per acre})_{j,y}}{(\text{surface area per tree})_{j,y}},(\text{TPA})_{j,y}\right),$ (8)
which clearly fits the data (Fig. 12). The slope parameter $\gamma$ has units $m^{2}/\text{attacks}$, so $1/\gamma$ is an estimate of attack density. The maximum number of killed trees is capped by the pre-attack trees per acre, $(\text{TPA})_{j,y}$.
The expected beetles in the following year is obtained by assuming that emergence density (per $\text{m}^{2}$) is a piecewise linear function of tree diameter, and then by summing across all $J$ diameter classes:
@@
$\mathbb{E}\left[B_{y+1}\right]=\sum_{j=1}^{J}\begin{cases}\overbrace{\left(\zeta_{0}+\zeta_{1}\cdot\text{DBH}_{j}\right)}^{\text{Emergence density}}(\overbrace{(\text{KTPA})_{j,y}\cdot(\text{surface area per tree})_{j}}^{\text{surface area of killed trees}}),&\text{if }\text{DBH}_{j}>-\frac{\zeta_{0}}{\zeta_{1}}\\ 0,&\text{otherwise}.\end{cases}$ (9)
Here, $\zeta_{0}$ and $\zeta_{1}$ are the intercept and slope parameters of the emergence $\sim$ DBH regression. The quantity $-\zeta_{0}/\zeta_{1}$ is the beetle-production threshold, i.e., trees with $\text{DBH}<-\zeta_{0}/\zeta_{1}$ will not produce any beetles, even if they are successfully attacked.
Figure 12: The number of killed trees is a linear function of successful beetle attacks (corrected for DBH-specific surface area), suggesting that beetles “fill up” a tree to a universal density, and then move onto the next tree. The prediction was generated using posterior means of model parameters. Data from Klein et al., (1978).
We cannot justify the functional form of Eq. 9 with graphical evidence, since the Klein et al., dataset only contains total emergence holes (not stratified by diameter class). However, the linear nature of the relationship has been well-established (Reid,, 1963, Table 1; Safranyik,, 1968, Fig. 25; Cole,, 1969, Fig. 5; Cudmore et al.,, 2009, Fig. 6). The linear relationship also implicitly accounts for the fact that the male:female sex ratio of emerging beetles increases with tree diameter (Cole et al.,, 1976, Fig. 5). Note that emergence only depends on killed trees, not attacking beetles, because a super-optimal attack density is balanced by a decrease in per capita oviposition or early-instar survival; this is perfectly compensatory density dependence at the tree scale (Fig. D.2; Berryman,, 1974, Fig. 2c; Berryman,, 1976, Fig. 2).
@@
To capture temporal variation in overwintering mortality, we assume that the realized number of beetles is drawn from a zero-truncated normal distribution:
@@
$B_{y+1}\sim\text{Normal}_{\scriptscriptstyle{Trunc.}}\left(\mathbb{E}\left[B_{y+1}\right],\sigma_{\text{env}}\mathbb{E}\left[B_{y+1}\right]\right).$ (10)
@@
2.3.2 Lodgepole pine dynamics
Each time-step of the simulation model begins and ends with the beetle flight period in July or August. Therefore, the first process affecting the lodgepole pine population is MPB-induced mortality. The density of remaining trees in late summer, denoted $(\text{TPA}^{\prime})_{j,y}$, is
$(\text{TPA}^{\prime})_{j,y}=(\text{TPA})_{j,y}-(\text{KTPA})_{j,y},\quad j=1,\ldots,J.$ (11)
Recall that $(\text{TPA})_{j,y}$ is the initial number of trees per acre, $(\text{KTPA})_{j,y}$ is the number of MPB-killed trees per acre, and $j=1,\ldots,J$ indexes the tree diameter class. The $J^{\text{th}}$ diameter class represents a mature individual that will not grow any larger.
After MPB-induced mortality, trees may die of auxiliary causes. This density-independent mortality is mostly likely to occur from late summer (drought stress) to late winter (ice damage and windthrow). The remaining $(1-m)$ proportion trees grow throughout the spring and summer, automatically entering the next diameter class. Now, the tree density is
$(\text{TPA})_{j+1,y+1}=(1-m)\times(\text{TPA}^{\prime})_{j,y}-(\text{KTPA})_{j,y},\quad j=2,\ldots,J-2.$ (12)
The $J^{\text{th}}$ diameter class receives individuals from the second-largest diameter class, but also retains the mature individuals:
$(\text{TPA})_{J,y+1}=(1-m)\times\left((\text{TPA}^{\prime})_{J-1,y}+(\text{TPA}^{\prime})_{J,y}\right).$ (13)
Trees are recruited into the smallest diameter class, such that there are always $K$ standing trees per acre, which includes live trees and MPB-killed trees that have not yet dropped their needles. We assume that MPB-killed trees do not create light gaps for 3 years, and that all density-independent mortality results in an immediate light gap. With these assumptions, recruitment can be written as
$(\text{TPA})_{1,y+1}=\overbrace{K}^{\text{Carrying capacity}}-\overbrace{\left(\sum_{j=1}^{J}\left(1-m\right)\times(\text{TPA}^{\prime})_{j,y}\right)}^{\begin{subarray}{c}\text{Trees remaining after}\\ \text{density-independent mortality}\end{subarray}}+\overbrace{\left(\sum_{s=y-2}^{y}\sum_{j=1}^{J}(\text{KTPA})_{j,s}\right)}^{\begin{subarray}{c}\text{MPB-killed snags}\\ \text{that still have needles}\end{subarray}}.$ (14)
@@ 2.3.2 Lodgepole pine dynamics
We assume that each tree’s DBH grows $4\text{mm}=0.157\text{in}$ each year, which is consistent with empirical estimates (Mata,, 2003; Heath and Alfaro,, 1990; McLane et al.,, 2011). However, like all pines, lodgepole pine grows faster in youth and slower in age, so 4mm here is a convenient approximation, representing the average growth increment over a tree’s lifespan.
Although lodgepole pine trees are capable of indeterminate growth, intraspecific competition in dense stands constrains the maximum diameter. The vast majority of lodgepole pine trees in even-aged stands do not exceed 20–24 inches DBH (Rodman et al.,, 2022; Negrón,, 2020; Huang,, 1999), and many U.S. Forest Service reports use a “$>18$ inches” category to describe lodgepole diameters (e.g., Cole,, 1969; Stipe,, 1974; McGregor,, 1973). Therefore, we assume that the trees stop growing after reaching 20 inches DBH. To summarize the fixed-value parameters associated with tree growth, we have $DBH_{1}=0.157$, $DBH_{J}=20-\left(20\bmod 0.157\right)\approx 20$, and $DBH_{j}=DBH_{j-1}+0.157$ for $j=2,\ldots,J$. Assuming that there are no MPB, and that $m$ is large enough that the $J^{th}$ DBH class contains very few individuals, we to can perform recursion on Eq. 12 to approximate the equilibrium distribution of tree diameters as $(\text{TPA})_{j,\infty}=(1-m)^{(j-1)}mK$. This geometric distribution matches the pre-outbreak distribution in the Klein et al., dataset (Fig. 13).
Figure 13: The pre-outbreak diameter distribution of lodgepole pine is effectively modelled by assuming density-independent mortality and instant recruitment (see 12 & Eq. 14). The theoretical distribution of trees per acre in the $j^{th}$ diameter class is $(\text{TPA})_{j,\infty}=(1-m)^{(j-1)}mK$, so the logarithm of trees per acre within should be approximately linear with respect to diameter at breast height (DBH). Data from Klein et al., (1978). The predictions were generated with posterior means of model parameters.
2.4 Model-fitting
The parameters of the simulation model are estimated with three statistical models: the lodgepole-dynamics model, the Allee-effect model, and the outbreak-dynamics model. The combined statistical models are conceptually identical to the simulation model, but their mathematical descriptions are more verbose, and thus they are relegated to Appendix A. All models were fit to data with Stan (Stan Development Team,, 2020), a Bayesian model-fitting software that implements Hamiltonian Monte Carlo.
[…]

### v2

@@ 2.1 Overview
To understand how MPB life history affects MPB population dynamics, we will build a mechanistic model of MPB dynamics (summarized in Fig. 5), then simulate the model under counterfactual scenarios where various biological features have been turned on or off. If a population-dynamical phenomenon (e.g., episodic outbreaks) disappears when a certain feature is turned off (e.g., stand-dependent dispersal mortality), we can say that the feature is a necessary condition for the phenomenon. If two features being turned off makes the difference, then the two features are jointly necessary. As always, our inferences are conditional on the approximate veracity of the model.
Figure 5: Depiction of the beetle/forest simulation model, and of the sub-models therein.
“Our model” is actually constituted by two types of related models: a simulation model and a set of statistical models. The statistical models are used for model-building and for parameter estimation. The parameter estimates feed into the simulation model, which is then used to explore system behavior under various scenarios. The description of the statistical model is quite long, due to a large number of sub-models, equations describing error distributions, and the extensive use of graphical evidence to justify model structure. Therefore, the statistical models are described in Appendix A, while the simulation model is described below (in Section 2.3).
Realism is the key feature of our model — we intend to emulate the actual ecological system so that we can conduct virtual experiments. Many models of MPB have been published, but many are too simple for our purposes, having favored mathematical tractability over realism. Other models are too complex, containing functional forms that cannot be reasonably parameterized given available data. For further discussion of previous MPB models, see the Discussion and Appendix B.5. In contrast, our models are thoroughly justified via clear patterns in the data and/or a preponderance of evidence from previous research (see Appendices A & B). All models were fit to data with Stan (Stan Development Team,, 2020), a Bayesian model-fitting software that implements Hamiltonian Monte Carlo (details in Appendix C).
@@ 2.2 Data
The model was parameterized using data from Boone et al., (2011) and Klein et al., (1978). All data was extracted from figures and tables using the GraphGrabber software and Microsoft Excel’s digitizing features. Specific variables are described in Table LABEL:tab:model_notation. The Boone et al., dataset contains the number of mass-attacked trees and the number of successfully mass-attacked trees, across a range of attack years, stands, and attack densities. This data justifies a functional form for the Allee effect. The second dataset, courtesy of Klein et al., (Fig. A.1), is one of the only datasets that contain beetle attacks, beetle emergence, and tree deaths, all stratified by year and tree diameter class. Our statistical model of MPB outbreak dynamics is primarily calibrated with this dataset, with the notable exception being that a Bayesian prior for the Allee effect parameter is derived from the Boone et al., data.
2.3 Simulation model
Mountain pine beetle outbreak dynamics (fast):
  1. 1.
    Beetle survival during the dispersal phase depends on forest conditions. The survival probability is a logistically increasing function of the total surface area of susceptible trees:
    $s_{y}=s_{0}\left(1+\exp\left(-\lambda_{1}\left((\text{relative surface area})_{y}-\lambda_{0}\right)\right)\right)^{-1}$ (1)
    where $s_{y}$ is the survival probability in the year $y$, and $s_{0}$ is the maximum survival probability. The parameters $\lambda_{0}$ $\lambda_{1}$ respectively control the intercept and slope of the logistic function. The relative surface area is the ratio of the current total surface area of trees with a diameter at breast height (DBH) of 8 inches or greater to the maximum possible surface area for such trees in an equilibrium forest without mountain pine beetle outbreaks.
  2. 2.
@@ 2.3 Simulation model
    Beetle populations experience stochastic perturbations. To obtain the density of attacking beetles, we multiply the number of emerging females per acre, $B_{y}$, and the dispersal-phase survival probability, $s_{y}$. Potentially, with probability $P_{I}$, we also add $I$ beetles per acre:
@@
    The Bernoulli random variable simulates the stochastic endemic $\rightarrow$ incipient-epidemic transition. Because most outbreaks occur during regional epidemics, beetle immigration is the proximal catalyst for many endemic $\rightarrow$ incipient-epidemic transitions on the local scale. But because drought is responsible for region-wide outbreaks, drought is the ultimate catalyst for most local-scale outbreaks. Drought conditions are somewhat unpredictable, but the best predictor of drought in western North America is the Pacific Decadal Oscillation (McCabe et al.,, 2004; Hidalgo,, 2004), which has a quasi-period of 15-30 years (Mantua and Hare,, 2002). The last two drought periods preceded the last two continent-scale outbreaks and were 20 years apart, (Creeden et al.,, 2014, Fig. 5d). Therefore, we set $P_{I}=1/20=0.05$. We also set $I=300$ based on a back-of-the-envelope calculation (Appendix B.1).
  3. 3.
    The probability of a successful mass attack is an exponential saturating function of attack density (i.e., beetles per acre). The density of successful attacks is
[…]
    where $a$ is the Allee threshold parameter (defined as the attack density where the success probability equals 0.9). Unsuccessful beetles die without attacking additional trees.
  4. 4.
    Beetles prefer larger trees. Beetle preference is a piecewise linear function of tree diameter where “preference” is operationalized as the proportion of beetle attacks within a particular diameter class, divided by the proportion of total tree surface area belonging to that class:
    $(\text{preference})_{j}=\begin{cases}\beta_{0}+\beta_{1}\cdot\text{DBH}_{j}&\text{if }\text{DBH}_{j}>-\beta_{0}/\beta_{1}\\ 0&\text{otherwise}.\end{cases}$ (4)
    Here, $\beta_{0}$ and $\beta_{1}$ are intercept and slope parameters; and $\text{DBH}_{j}$ is the diameter at breast height of the $j^{\text{th}}$ discrete diameter class. Thus, $-\beta_{0}/\beta_{1}$ is the susceptible DBH threshold for beetle attacks.
  5. 5.
    Beetles distribute among tree diameter classes based on preference. The number of beetles per acre that successfully attack the $j^{th}$ DBH class is:
@@
  6. 6.
    A tree is killed once it is “filled-up” to a density of $1/\gamma$ female beetles per $\text{m}^{2}$. The number of killed trees per acre (KTPA) in each diameter class is:
@@
    $(\text{KTPA})_{j,y}=\text{min}\left(\gamma\frac{(\text{succ. attacks per acre})_{j,y}}{(\text{surface area per tree})_{j,y}},(\text{TPA})_{j,y}\right)$ (6)
    where $\gamma$ a slope parameter, and $(\text{TPA})_{j,y}$ is trees per acre. If “successfully” attacking beetles kill all of the trees within a size class, the excess beetles are assumed to die (hence the “min” function).
  7. 7.
    The number of beetles produced by a single tree increases quadratically with tree diameter. Beetle emergence density (emerging female beetles per $\text{m}^{2}$) is a piecewise linear function of tree DBH:
@@
    $\left(\text{emerge density}\right)_{j,y}=\begin{cases}\zeta_{0}+\zeta_{1}\cdot\text{DBH}_{j}&\text{if }\text{DBH}_{j}>-\zeta_{0}/\zeta_{1}\\ 0&\text{otherwise}\end{cases}$ (7)
    Here, $\zeta_{0}$ and $\zeta_{1}$ are intercept and slope parameters. Thus, $-\zeta_{0}/\zeta_{1}$ is the DBH threshold for beetle reproduction. The expected number of female beetles per acre in next year, $y+1$, is obtained with simple accounting:
    $\mathbb{E}\left[B_{y+1}\right]=\sum_{j=1}^{J}\left(\text{emerge density}\right)_{j,y}\cdot(\text{surface area per tree})_{j}\cdot(\text{KTPA})_{j,y}.$ (8)
  8. 8.
@@
    The realized number of beetles is stochastic. To capture temporal variation in overwintering mortality, we assume that the realized number of beetles is drawn from a zero-truncated normal distribution:
@@
    $B_{y+1}\sim\text{Normal}_{\scriptscriptstyle{Trunc.}}\left(\mathbb{E}\left[B_{y+1}\right],\sigma_{\text{env}}\cdot\mathbb{E}\left[B_{y+1}\right]\right).$ (9)
@@
Lodgepole pine dynamics (slow):
  1. 1.
    Trees grow at a constant rate until maturity. Tree growth is 4 mm (0.157 inches) DBH increase per year until trees reach 20 inches DBH, after which growth stops.
@@
    Although lodgepole pine trees are capable of indeterminate growth, intraspecific competition in dense stands constrains the maximum diameter. The vast majority of lodgepole pine trees in even-aged stands do not exceed 20–24 inches DBH (Rodman et al.,, 2022; Negrón,, 2020; Huang,, 1999), and many U.S. Forest Service reports use a “$>18$ inches” category to describe lodgepole diameters (e.g., Cole,, 1969; Stipe,, 1974; McGregor,, 1973). The DBH growth rate is consistent with empirical estimates (Mata,, 2003; Heath and Alfaro,, 1990; McLane et al.,, 2011). However, like all pines, lodgepole pine grows faster in youth and slower in age, so 4mm here is a convenient approximation, representing the average increment over a tree’s lifespan.
  2. 2.
    Trees experience constant background mortality. A fixed proportion $(m)$ of trees die each year, regardless of size:
    $(\text{TPA})_{j+1,y+1}=(1-m)\times(\text{TPA}^{\prime})_{j,y}-(\text{KTPA})_{j,y}$ (10)
    where $(\text{TPA})_{j,y}$ is trees per acre in diameter class $j$ and year $y$, and $(\text{TPA}^{\prime})_{j,y}$ is trees per acre remaining after beetle-induced mortality.
  3. 3.
    Forest maintains constant density through recruitment. Recruitment into the smallest diameter class (0.157 inches DBH) ensures a $K$ standing trees per acre, which includes live trees and MPB-killed trees that have not yet dropped their needles:
    $(\text{TPA})_{1,y+1}=K-\sum_{j=1}^{J}\left(1-m\right)\times(\text{TPA}^{\prime})_{j,y}+\sum_{s=y-2}^{y}\sum_{j=1}^{J}(\text{KTPA})_{j,s}.$ (11)
  4. 4.
    Dead trees temporarily occupy canopy space. Trees killed by MPB retain needles for 3 years, occupying forest canopy and supressing recruitment as a live tree would.